can you re write these problems by using distributive property
7(2x-3)
-3 (2x-5)
8x (-3x-7)
And can you undistribute these problems
6x+18

Answer by BurningFlame_marc(77)

Hello!
To distribute you will multiply the number outside the parentheses to the numbers inside the parentheses.
7(2x-3) = 14x-21
-3 (2x-5) = -6x+15
8x (-3x-7) = -24x^2 -56x

To undistribute you will divide the numbers by their common factor then put their common factor outside the parenthese and put the answers inside.

6x+18 = 6(x+3)
Hope you learned from me!